# Branch length statistics in phylogenetic trees under constant-rate birth-death dynamics

### v3

@@ VIII Discussion
For the first scenario, we generalize the approach of Paradis Paradis (2016) of observing the population at a given time to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic distribution. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. An interesting consequence of sampling at a low probability concerns the number of generations (birth events which have not lead to additional sampled clades) along interior branches: even at low sampling probabilities, the number of generations along interior branches can be small due to the short branch lengths. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than by a continuous accumulation over time. The consequences and applications of this result will be explored in a future publication.
@@
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (such as a branch of a given length), relative to the probability of a population with a certain size. The combination of first going backwards and then forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). The ”going backwards in time”-aspect of this approach is shared with coalescent approaches, like the Kingman-coalescent Kingman (1982). However, the statistics we compute follows that of the original birth-death process which generated the tree. It thus includes the effects of lineages that died out (reflected for instance in the number of generations along a branch), and the distinct behaviour of pendant and interior branches under sampling, which are generally missing in coalescent approaches. Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches.
@@
The interior branch distributions of the asymptotic cases of both scenarios (large times for scenario i) and large population sizes for ii) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size $N$, even when $N$ is not large, as the latter distribution is independent of $N$ Stadler and Steel (2012).
@@ Acknowledgements.
$p_{N,n}^{\lambda,\mu}(t)$ $=\sum_{j=0}^{\min(n,N)}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}\alpha^{n-j}\beta^{N-j}(1-\alpha-\beta)^{j}$ (25)
with shorthand
$\alpha=\alpha(t,\lambda,\mu)$ $=\mu\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-1}{\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-\mu}$ (26)
$\beta=\beta(t,\lambda,\mu)$ $=\lambda\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-1}{\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-\mu}\ .$ (27)
This probability can be derived by considering $n$ clades starting out with a single individual at a particular time and evolving according to (4), and computing the probability that the sum of all clade sizes a time $t$ later is $N$. For $N=0$ equation (25) simplifies to $p_{0,n}^{\lambda,\mu}(t)=\alpha^{n}=p_{0}(\tau)^{n}$.
The conditional probability (25) can also be used to describe that statistics of the population size $n$ some time $\tau$ in the past, conditioned on a population size $N$ at the time of observation. Ignatieva, Hein, and Jenkins Ignatieva et al. [2020] show that the statistics of $n$ is described by a birth-death process running for time $\tau$, starting with $N$ and with the rates of birth and death exchanged. This result uses a flat prior on the (unknown) age $T$ of the birth-death process. The same flat prior is also used explicitly in the derivation of the pendant branch length distribution conditioned on the final population size by Stadler and collaborators Mooers et al. [2012], Stadler and Steel [2012].
An important observation that simplifies the subsequent calculations is that the ratio of conditional probabilities going forwards and backwards in time simplifies to
$\frac{p_{n,N}^{\mu,\lambda}(\tau)}{p_{N,n}^{\lambda,\mu}(\tau)}=N/n\ .$ (28)
for all times and rates. This can be derived using the symmetry in the exchange of birth and death rates $\alpha(t,\mu,\lambda)=\beta(t,\lambda,\mu)$. Ignoring the binomial coefficients in equation (25), the remaining terms are therefore invariant to swapping $\lambda\leftrightarrow\mu$ and $N\leftrightarrow n$. For the binomial coefficients we get (${{N}\choose{j}}{{N+n-j-1}\choose{N-1}}=\frac{N}{n}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}$. Since the upper limit of the sum in equation (25) is invariant under exchange of $n$ and $N$, we get the result (28).
@@
We now solve the toy problem of Section VII (pendant branch length distribution conditioned on a final population of size $N=2$) in three different ways of increasing generalizability. The last one generalizes directly to the result for any value of the final population size $N$ and to the goal of this appendix, the distribution of interior branch lengths conditioned on population size.
The first approach takes the sum (24), which with (28) becomes the straightforward geometric sum
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{n,2}^{\mu,\lambda}(\tau)}{p_{2,n}^{\lambda,\mu}(\tau)}=2\lambda p_{1}(\tau)^{2}\sum_{n=1}^{\infty}p_{0}(\tau)^{n-1}=2\lambda\frac{p_{1}(\tau)^{2}}{1-p_{0}(\tau)}\ .$ (29)
In the second approach, we look at all individuals just after the cell division event at time $\tau$ that establishes the pendant branch we are considering. Figure 5A sketches the situation: There are $n+1$ individuals, one of which has just established the pendant branch. This individual will have exactly one extant offspring at the time of observation. The other $n$ individuals must also have exactly one extant offspring in total, leading to a population size of $N=2$ at the time of observation. Additionally, the particular individual that was also born at time $\tau$ (but does not establish the branch we consider) must have at least one extant offspring. (For the case $N=2$, the branches starting with the birth event at time $\tau$ are equivalent, but these statements generalize to $N>2$). The probability of the $n$ individuals having one extant offspring at the time of observation, with one given individual not going extinct is $P(1|n)-p_{0}P(1|n-1)$ with shorthand $P(N|n)\equiv p_{N,n}^{\lambda,\mu}(\tau)$, $p_{0}=p_{0}(\tau)$. (Similarly, we do not note all dependencies of $\tau$ in the following equations, i.e. $p_{1}=p_{1}(\tau)$.) In this approach, we rewrite the relative fraction of birth events leading to a pendant branch at time $\tau$ (24) as
$\sum_{n=1}^{\infty}\lambda np_{1}(P(1|n)-p_{0}P(1|n-1))\frac{p_{n,2}^{\mu,\lambda}}{p_{2,n}^{\lambda,\mu}}=2\lambda p_{1}\sum_{n=1}^{\infty}(P(1|n)-p_{0}P(1|n-1))=2\lambda p_{1}(1-p_{0})\sum_{n=1}^{\infty}P(1|n)\ ,$ (30)
which gives the same result as above since $P(1|n)=np_{1}p_{0}^{n-1}$ and $\sum_{n=1}^{\infty}np_{1}p_{0}^{n-1}=p_{1}\partial_{p_{0}}\frac{p_{0}}{1-p_{0}}=\frac{p_{1}}{(1-p_{0})^{2}}$.
@@
Figure 5: Formation of branches given population size $N$ (A) Given a population size $N$ at the time of observation and $n$ individuals at time $\tau$ in the past, a pendant branch (red) forms from a birth event where one offspring has exactly one and the other offspring has at least one surviving descendant. The number of descendants from the latter and the other $n-1$ individuals at time $\tau$ must equal $N-1$. (B) The same situation, but for an interior branch: At time $\tau_{s}$ one of the $n$ individuals undergoes a birth event. To form an interior branch (red), one offspring has to form a lineage until time $\tau_{e}$ where all additional lineages arising die out before the population is observed. At time $\tau_{e}$, the individual undergoes a birth event with both offspring forming observed clades with a total of $m\geq 2$ individuals. The second offspring from the birth event at $\tau_{s}$ cannot die out but forms a clade of size $r\geq 1$. The remaining $N-m-r$ observed individuals have to be offspring of the $n-1$ individuals that did not divide at time $\tau_{s}$.
The third approach uses a generating function to compute $P(1|n)$. The probability that a given clade has size $n$ at a time $\tau$ after it originated with a single individual is given by (2)-(4) and can be expressed as
$p_{n}(\tau)$ $=p_{n,1}^{\lambda,\mu}(\tau)=(1-\alpha)(1-\beta)\beta^{n-1}=p_{1}(\tau)\left(\frac{\lambda}{\mu}p_{0}(\tau)\right)^{n-1}$ $n$ $>0$
$p_{0}(\tau)$ $=p_{0,1}^{\lambda,\mu}(\tau)=\alpha$ $n$ $=0\ .$ (31)
@@
$g(x)=\sum_{k=0}^{\infty}p_{k}(\tau)x^{k}=\alpha+(1-\alpha)(1-\beta)\sum_{k=1}^{\infty}\beta^{k-1}x^{k}\ .$ (32)
$P(1|n)$ is the coefficient of $x$ to the first power in the $n$th power of the generating function $g(x)$, which is $P(1|n)=\partial_{x}|_{x=0}g(x)^{n}=np_{1}p_{0}^{n-1}$ as in the second approach. However, since the geometric sum over $g(x)^{n}$ converges absolutely ($|\alpha|<1$ and $|\beta|<1$ for $\mu<\lambda$ and $\tau>0$), the summation over $n$ and differentiation commute, so we can perform the sum over $n$ first
@@
$\frac{g(x)}{1-g(x)}=\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\ .$ (34)
@@
$\sum_{n=1}^{\infty}P(1|n)=\partial_{x}|_{x=0}\left[\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\right]=\frac{1-\beta}{1-\alpha}=p_{1}/(1-p_{0})^{2}$ (35)
@@
The third approach based on the generating function generalizes immediately to any given final population size. With (28) the relative fraction of birth events leading to a pendant branch is instead of (30)
@@
$\sum_{n=1}^{\infty}P(N-1|n)=\frac{1}{(N-1)!}\partial_{x}^{N-1}|_{x=0}\left[\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\right]=p_{1}/(1-p_{0})^{2}\ ,$ (37)
independently of the final population size $N$. The resulting normalization factor turns out to be $2/N$, which means that the distribution of pendant branch lengths conditioned on the population size $N$ is independent of $N$.
Superficially, the result (29) looks different from the distribution of pendant branch lengths derived in Stadler and Steel Stadler and Steel [2012] and Mooers et al. Mooers et al. [2012], which in the present notation is $2\lambda p_{1}(\tau)\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)$. The results agree, as they must, since $\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)=(1-\beta)=\frac{p_{1}(\tau)}{1-p_{0}(\tau)}$ (compare to (.1)). Furthermore, the result coincides with the distribution of pendant branches conditioned on large tree ages $T$ (10) derived in Section III, as $p_{1}(\tau)/(1-p_{0}(\tau))=\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}(1-p_{0}(\tau))$.
@@
The approach based on generating functions can also be applied straightforwardly to the statistics of interior branches conditioned on the final population size $N$. Figure 5B shows how a population of size $n$ at time $\tau_{s}$ gives rise to an interior branch. The branch we focus on must have $m\geq 2$ extant offspring at the time of observation (with only one offspring it would be a pendant branch), the second individual born at time $\tau_{s}$ must have $r\geq 1$ offspring, and the remaining $n-1$ individuals have in total $N-r-m$ offspring. The latter condition ensures that the final population size equals $N$. The probability for such a trajectory starting from $n+1$ individuals at time $\tau_{s}$ and ending with $N$ at the time of observation is
@@
The next step looks at the clade establishing the interior branch at time $\tau_{s}$. Given that this clade has size $m\geq 2$ at the time of observation, what is the probability density (per small time period) of a birth event at time $\tau_{e}$, while the branch has no offspring with extant descendants prior to $\tau_{e}$ but is terminated at $\tau_{e}$, because both offspring establish extant lineages. This is
@@
$\frac{p_{1}^{\tau_{e}}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}}{1-p_{0}(\tau_{s})-p_{1}(\tau_{s})}\sum_{n=1}^{\infty}\lambda n\frac{p_{n,N}^{\mu,\lambda}(\tau_{s})}{p_{N,n}^{\lambda,\mu}(\tau_{s})}p_{\mbox{mr}(\tau_{s})}\ .$ (40)
@@
In the case of pendant branches, this means we do not know the exact number of extant descendants left by the $n$ individuals after the birth event at time $\tau$ who do not form the pendant branch (in the case of complete sampling they need to be $N-1$). However, since equation (37) is independent of $N$, summing over all possible numbers of descendants leads to the same factor, and the density function of pendant branches is thus (up to normalization) given by equation (29) with $p_{0}^{(\rho)}(\tau)$ and $p_{1}^{(\rho)}(\tau)$, coinciding with the result for trees at a given large age $T$ of the birth-death process in the case of finite sampling (18).

### v4

@@ VIII Discussion
For the first scenario, we generalize the approach of Paradis Paradis (2016) of observing the population at a given time to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic distribution. An intuitive explanation of this effect is given in Section VI. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. An interesting consequence of sampling at a low probability concerns the number of generations along interior branches (birth events which have not led to additional sampled clades): even at low sampling probabilities, the number of generations along interior branches can be small due to the short branch lengths. In contrast to macroevolution, where mutations arise at some rate per unit time, under cellular reproduction mutations can also occur at discrete moments in time at birth events Spisak et al. (2023). This means that the number of mutations along a branch can be shaped by the (few) birth events along that branch, which affects the statistics of mutations and leads to a compound Poisson statistics replacing the Poisson statistics that typically arises from the continuous accumulation of mutations. The interplay between reproduction and mutations also affects the statistics of mutations along a lineage Cheek and Johnston (2023). In Dieselhorst and Berg (2024), we use the mutation statistics along branches to infer the parameters of the underlying birth-death model.
@@
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (such as a branch of a given length), relative to the probability of a population with a certain size. The combination of first going backwards and then forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). The “going backwards in time”-aspect of this approach is shared with coalescent approaches, like the Kingman-coalescent Kingman (1982). However, the statistics we compute follows that of the original birth-death process which generated the tree. It thus includes the effects of lineages that died out, and the distinct behaviour of pendant and interior branches under sampling, which are generally missing in coalescent approaches.
@@
Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches. The interior branch distributions of the asymptotic cases of both scenarios (large times for scenario i) and large population sizes for ii)) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size $N$, even when $N$ is not large.
These results all hinge upon the assumption of a constant-rate birth-death model. Changes in birth and death rates can occur over time or between different clades, and this would affect the statistics of branch lengths. In turn, such changes in the statistics of branch lengths allow one in principle to infer rate changes from data Dieselhorst and Berg (2024). Also, in a cellular setting, the rates of birth depend on the cell cycle Mulberry et al. (2025). Calculating the resulting branch length distributions remains a task to be undertaken in future work.
@@ Acknowledgements.
${p_{(N,n)}^{(\lambda,\mu)}(t)}$ $=\sum_{j=0}^{\min(n,N)}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}{p_{0}^{n-j}\left(\frac{\lambda}{\mu}p_{0}\right)^{N-j}(1-p_{0}-\frac{\lambda}{\mu}p_{0})^{j}}$ (25)
with shorthand $p_{0}=p_{0}(t)=p_{0}(t,\lambda,\mu)$ (equation (2)). (The explicit dependence of $p_{0}$ on $\lambda$ and $\mu$ will become relevant later on.) The probability (25) can be derived by considering $n$ clades starting out with a single individual at a particular time and evolving according to (4), and computing the probability that the sum of all clade sizes a time $t$ later is $N$. For $N=0$ equation (25) simplifies to $p_{(0,n)}^{(\lambda,\mu)}(t)=p_{0}(\tau)^{n}$.
The conditional probability (25) can also be used to describe the statistics of the population size $n$ some time $\tau$ in the past, conditioned on a population size $N$ at the time of observation. Ignatieva, Hein, and Jenkins Ignatieva et al. [2020] show that the statistics of $n$ is described by a birth-death process running for time $\tau$, starting with $N$ and with the rates of birth and death exchanged. This result uses a flat prior on the (unknown) age $T$ of the birth-death process. The same flat prior is also used explicitly in the derivation of the pendant branch length distribution conditioned on the final population size by Stadler and collaborators Mooers et al. [2012], Stadler and Steel [2012].
An important observation that simplifies the subsequent calculations is that the ratio of conditional probabilities going forwards (from $n$ to $N$ individuals with birth rate $\lambda$ and death rate $\mu$) and backwards (from $N$ to $n$ individuals with birth rate $\mu$ and death rate $\lambda$) in time simplifies to
${\frac{p_{(n,N)}^{(\mu,\lambda)}(\tau)}{p_{(N,n)}^{(\lambda,\mu)}(\tau)}}=N/n\ .$ (26)
for all times and rates. This can be derived using the symmetry in the exchange of birth and death rates $p_{0}(t,\mu,\lambda)=\frac{\lambda}{\mu}p_{0}(t,\lambda,\mu)$. Ignoring the binomial coefficients in equation (25), the remaining terms are therefore invariant to swapping $\lambda\leftrightarrow\mu$ and $N\leftrightarrow n$. For the binomial coefficients we get ${{N}\choose{j}}{{N+n-j-1}\choose{N-1}}=\frac{N}{n}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}$. Since the upper limit of the sum in equation (25) is invariant under exchange of $n$ and $N$, we get the result (26).
@@
We start by solving the toy problem of Section VII (pendant branch length distribution conditioned on a final population of size $N=2$) in two alternative ways of increasing generalizability. The second one generalizes directly to the result for any value of the final population size $N$ and to the goal of this appendix, the distribution of interior branch lengths conditioned on population size. In the following, we will use the shorthand $P(N|n)\equiv{p_{(N,n)}^{(\lambda,\mu)}(\tau)}$ and do not denote all dependencies on $\tau$, i.e. $p_{0}=p_{0}(\tau)$ and $p_{1}=p_{1}(\tau)$.
In the first approach, we look at all individuals just after the cell division event at time $\tau$ that establishes the pendant branch we are considering. Figure 5A sketches the situation: There are $n+1$ individuals, one of which has just established the pendant branch. This individual will have exactly one extant offspring at the time of observation. The other $n$ individuals must also have exactly one extant offspring in total (which happens with probability $P(1\mid n)$), leading to a population size of $N=2$ at the time of observation. Additionally, the particular individual that was also born at time $\tau$ (but does not establish the branch we consider) must have at least one extant offspring. (For the case $N=2$, the branches starting with the birth event at time $\tau$ are equivalent, but these statements generalize to $N>2$). The probability of the $n$ individuals having one extant offspring at the time of observation, with one given individual not going extinct is $P(1|n)-p_{0}P(1|n-1)$, where the latter term excludes the case where the $n$ individuals leave the desired single descendant, but the lineage of the second offspring of the birth event forming the pendant branch has died out. In this approach, we rewrite the relative fraction of birth events leading to a pendant branch at time $\tau$ (left-hand side of (24)) as
$\sum_{n=1}^{\infty}\lambda np_{1}\left(P(1|n)-p_{0}P(1|n-1)\right){\frac{p_{(n,2)}^{(\mu,\lambda)}}{p_{(2,n)}^{(\lambda,\mu)}}}=2\lambda p_{1}\sum_{n=1}^{\infty}\left(P(1|n)-p_{0}P(1|n-1)\right)=2\lambda p_{1}(1-p_{0})\sum_{n=1}^{\infty}P(1|n)\ ,$ (27)
which gives the same result as before (equation (24)) since $P(1|n)=np_{1}p_{0}^{n-1}$ and $\sum_{n=1}^{\infty}np_{1}p_{0}^{n-1}=p_{1}\partial_{p_{0}}\frac{p_{0}}{1-p_{0}}=\frac{p_{1}}{(1-p_{0})^{2}}$.
@@
Figure 5: Formation of branches given population size NN (A) Given a population size $N$ at the time of observation and $n$ individuals at time $\tau$ in the past, a pendant branch (red) forms from a birth event where one offspring has exactly one and the other offspring has at least one surviving descendant. The number of descendants from the latter and the other n1n-1 individuals at time $\tau$ must equal N1N-1. (B) The same situation, but for an interior branch: At time $\tau_{s}$ one of the $n$ individuals undergoes a birth event. To form an interior branch (red), one offspring has to form a lineage until time $\tau_{e}$ where all additional lineages arising die out before the population is observed. At time $\tau_{e}$, the individual undergoes a birth event with both offspring forming observed clades with a total of $m\geq 2$ individuals. The second offspring from the birth event at $\tau_{s}$ cannot die out but forms a clade of size $r\geq 1$. The remaining NmrN-m-r observed individuals have to be offspring of the n1n-1 individuals that did not divide at time $\tau_{s}$.
The second approach uses a generating function to compute $P(1|n)$. A generating function $g(x)\equiv\sum_{n}p_{n}x^{n}$ represents a probability distribution $p_{n}$ over a discrete set of outcomes as the coefficients of a power series. The coefficient $p_{n}$ can be recovered by taking the $n$-th derivative of $g$ with respect to $x$. The generating function can be interpreted as the expectation value of $x^{n}$, i.e. $g(x)=\mathrm{E}\left[x^{n}\right]$.
The probability that a given clade has size $n$ at a time $\tau$ after it originated with a single individual is given by (2)-(4) and can be expressed as
$p_{n}(\tau)$ $={p_{(n,1)}^{(\lambda,\mu)}(\tau)=(1-p_{0})\left(1-\frac{\lambda}{\mu}p_{0}\right)\left(\frac{\lambda}{\mu}p_{0}\right)^{n-1}=p_{1}\left(\frac{\lambda}{\mu}p_{0}\right)^{n-1}}$ $n$ $>0$
$p_{0}(\tau)$ $={p_{(0,1)}^{(\lambda,\mu)}(\tau)=p_{0}}$ $n$ $=0\ .$ (28)
@@
$g(x)={\mathrm{E}\left[x^{n}\right]=}\sum_{k=0}^{\infty}p_{k}(\tau)x^{k}={p_{0}+(1-p_{0})\left(1-\frac{\lambda}{\mu}p_{0}\right)\sum_{k=1}^{\infty}\left(\frac{\lambda}{\mu}p_{0}\right)^{k-1}x^{k}}\ .$ (29)
This is an infinite geometric series. When starting with $n$ individuals, the total population size $N$ at time $\tau$ is described by the sum of the $n$ independent and identically distributed sizes $k_{i}$ of the clades generated by each individual, i.e. $N=\sum_{i=1}^{n}k_{i}$. Because $\mathrm{E}\left[x^{N}\right]=\mathrm{E}\left[x^{k}\right]^{N}$, the probability generating function of $N$ is simply the $n$th power of $g(x)$. Furthermore, $P(1|n)$ is the coefficient of $x$ to the first power in $g(x)^{n}$ and therefore $P(1|n)=\partial_{x}|_{x=0}g(x)^{n}=np_{1}p_{0}^{n-1}$ as in the second approach. However, since the geometric sum over $g(x)^{n}$ converges absolutely ($|{p_{0}}|<1$ and $\left|{\frac{\lambda}{\mu}p_{0}}\right|<1$ for $\mu<\lambda$ and $\tau>0$), the summation over $n$ and differentiation commute, so we can perform the sum over $n$ first
@@
$\frac{g(x)}{1-g(x)}={\frac{p_{0}}{1-p_{0}}\frac{1}{1-x}+\frac{1-p_{0}-\frac{\lambda}{\mu}p_{0}}{1-p_{0}}\frac{x}{1-x}}\ .$ (31)
@@
$\sum_{n=1}^{\infty}P(1|n)=\partial_{x}|_{x=0}{\left[\frac{p_{0}}{1-p_{0}}\frac{1}{1-x}+\frac{1-p_{0}-\frac{\lambda}{\mu}p_{0}}{1-p_{0}}\frac{x}{1-x}\right]=\frac{1-\frac{\lambda}{\mu}p_{0}}{1-p_{0}}=\frac{p_{1}}{(1-p_{0})^{2}}}$ (32)
@@
The second approach based on the generating function generalizes immediately to any given final population size. With (26) the relative fraction of birth events leading to a pendant branch is instead of (27)
@@
$\sum_{n=1}^{\infty}P(N-1|n)=\frac{1}{(N-1)!}\partial_{x}^{N-1}|_{x=0}{\left[\frac{p_{0}}{1-p_{0}}\frac{1}{1-x}+\frac{1-p_{0}-\frac{\lambda}{\mu}p_{0}}{1-p_{0}}\frac{x}{1-x}\right]=\frac{1-\frac{\lambda}{\mu}p_{0}}{1-p_{0}}=\frac{p_{1}}{(1-p_{0})^{2}}}\ ,$ (34)
independently of the final population size $N$. Here, we have again used that the coefficient of the (N1)(N-1)th power of $x$ in $g(x)^{n}$ yields $P(N-1\mid n)$. The resulting normalization factor turns out to be $2/N$, which means that the distribution of pendant branch lengths conditioned on the population size $N$ is independent of $N$ and thus given by (24).
Superficially, the result (24) looks different from the distribution of pendant branch lengths derived in Stadler and Steel Stadler and Steel [2012] and Mooers et al. Mooers et al. [2012], which in the present notation is $2\lambda p_{1}(\tau)\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)$. The results agree, as they must, since $\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)=\frac{p_{1}(\tau)}{1-p_{0}(\tau)}$ (compare to (.1)). Furthermore, the result coincides with the distribution of pendant branches conditioned on large tree ages $T$ (10) derived in Section III, as ${\frac{p_{1}(\tau)}{(1-p_{0}(\tau))}}=\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}(1-p_{0}(\tau))$.
@@
The approach based on generating functions can also be applied straightforwardly to the statistics of interior branches conditioned on the final population size $N$. Figure 5B shows how a population of size $n$ at time $\tau_{s}$ gives rise to an interior branch. The branch we focus on must have $m\geq 2$ extant offspring at the time of observation (with only one offspring it would be a pendant branch), the second individual born at time $\tau_{s}$ must have $r\geq 1$ offspring, and the remaining n1n-1 individuals have in total NrmN-r-m offspring. The latter condition ensures that the final population size equals $N$. The probability for such a trajectory starting from $n+1$ individuals at time $\tau_{s}$ and ending with $N$ at the time of observation is
@@
The next step looks at the clade establishing the interior branch at time $\tau_{s}$. Given that this clade has size $m\geq 2$ at the time of observation, what is the probability density (probability per small time period) of a birth event at time $\tau_{e}$, while the branch has no offspring with extant descendants prior to $\tau_{e}$ but is terminated at $\tau_{e}$, because both offspring establish extant lineages. This is
@@
$\frac{p_{1}^{\tau_{e}}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}}{1-p_{0}(\tau_{s})-p_{1}(\tau_{s})}\sum_{n=1}^{\infty}\lambda n{\frac{p_{(n,N)}^{(\mu,\lambda)}(\tau_{s})}{p_{(N,n)}^{(\lambda,\mu)}(\tau_{s})}}p_{\mbox{mr}(\tau_{s})}\ .$ (37)
@@
In the case of pendant branches, this means we do not know the exact number of extant descendants left by the $n$ individuals after the birth event at time $\tau$ who do not form the pendant branch (in the case of complete sampling they need to be N1N-1). However, since equation (34) is independent of $N$, summing over all possible numbers of descendants leads to the same factor, and the density function of pendant branches is thus (up to normalization) given by equation (24) with $p_{0}^{(\rho)}(\tau)$ and $p_{1}^{(\rho)}(\tau)$, coinciding with the result for trees at a given large age $T$ of the birth-death process in the case of finite sampling (18).